### 24 Mar 2026 v2

# Maximum packings in graphs forbidding given rainbow cycles

Abstract
Motivated by the Ruzsa-Szemerédi problem, Imolay, Karl, Nagy, and Váli studied a variant of Turán number $ex_{F}(n,G)$ (called the $F$-multicolor Turán number of $G$), defined as the maximum number of edge-disjoint copies of $F$ on $n$-vertex set such that there is no copies of $G$ whose edges come from distinct copies of $F$. They proved that if there is no homomorphism from $G$ to $F$, then $n^{2}/v(F)^{2}+o(n^{2})\leq ex_{F}(n,G)\leq ex(n,G)/e(F)+o(n^{2})$, and otherwise $ex_{F}(n,G)=o(n^{2})$. The quantity $ex_{F}(n,G)$ asymptotically equals the maximum size of an $F$-packing in an $n$-vertex $G$-free graph, and attains the upper bound $ex(n,G)/e(F)+o(n^{2})$ if and only if $\chi(G)>\chi(F)$.
In this paper, we provide conditions under which $ex_{F}(n,G)$ does not achieve the lower bound $n^{2}/v(F)^{2}+o(n^{2})$, and describe additional graph pairs that attain this lower bound via graph blow-ups. Especially, we proved that $ex_{C_{k}(s)}(n,C_{k-2})=n^{2}/(sk)^{2}+o(n^{2})$ for any $k\geq 5$. For degenerate cases, we show that if $\chi(F)=3$ and $G$ and $F$ share the same odd girth, then $ex_{F}(n,G)$ satisfies the $(6,3)$-type bound $n^{2-o(1)}$, generalizing a result of Kovács and Nagy. We also prove that $ex_{C_{2k+1}}(n,C_{2\ell+1})=O(n^{1+1/(\ell-k+1)})$ for any integers $k,\ell$ with $\ell>k$, extending a result of Füredi and Özkahya. Additionally, we establish $ex_{C_{4}}(n,C_{4})=\sqrt{2}n^{3/2}/8+O(n)$.
Keywords: $F$-multicolor Turán number, packing, cycles, the lower bound, the $(6,3)$-type bound
AMS subject classification 2020: 05C35.
1 Introduction
The $(6,3)$-problem can be equivalently stated as follows: determine the maximum number of edge-disjoint triangles in an $n$-vertex graph such that no triangle is formed by taking one edge from each of three distinct triangles. Imolay, Karl, Nagy and Váli [13] generalized this problem by considering the following: given two fixed graphs $F$ and $G$, determine the maximum number of edge-disjoint copies of $F$ in an $n$-vertex graph such that there is no copy of $G$ whose edges come from distinct copies of $F$. This maximum, denoted by $ex_{F}(n,G)$, is called the $F$-multicolor Turán number of $G$. Clearly, if $F$ and $G$ are $r$-uniform hypergraphs, then $ex_{F}(n,G)=O(n^{r})$.
For convenience, we may regard a collection of edge-disjoint copies of $F$ as a family of monochromatic copies of $F$, each assigned a distinct color. Any such edge-disjoint copy of $F$ is called a monochromatic $F$-copy, while any other (not necessarily monochromatic) copy of $F$ is simply called a copy of $F$. Then $ex_{F}(n,G)$ is the maximum number of monochromatic $F$-copies in an $n$-vertex graph that contains no rainbow copy of $G$. The $F$-multicolor Turán number is a very interesting concept, as it is closely related to three classical problems: the linear Turán number of Berge graphs, the packing number, and the generalized Turán number.
For a family $\mathcal{G}$ of graphs, the linear Turán number of Berge-$\mathcal{G}$, denoted $ex_{r}^{lin}(n,\mathcal{G})$, is the maximum number of edges in an $n$-vertex linear $r$-uniform hypergraph that contains no Berge copy of any $G\in\mathcal{G}$ (i.e., no copy of $G$ whose edges come from distinct hyperedges). If $F=K_{r}$ is a complete graph, then any two monochromatic $F$-copies intersect in at most one vertex. Regarding each such monochromatic $F$-copy as an $r$-edge, we obtain $ex_{F}(n,G)=ex_{r}^{lin}(n,G)$.
The $G$-packing number of a graph $H$, denoted $\nu_{H}(G)$, is the maximum number of edge-disjoint copies of $G$ in $H$. An interesting problem is to determine the packing number in graphs that forbid some given graphs. We use $\nu_{n}(F,G)$ to denote the maximum $F$-packing number among all $n$-vertex $G$-free graphs. It is worth noting that $ex_{F}(n,G)$ and $\nu_{n}(F,G)$ are deeply related.
Proposition 1.1.
For any two $r$-graphs $G$ and $F$, $\nu_{n}(F,G)=ex_{F}(n,G)+o(n^{r})$.
Proposition 1.1 holds because the inequality $\nu_{n}(F,G)\leq ex_{F}(n,G)$ is trivial, and conversely,
  • ($\star$).
    for an $n$-vertex $r$-graph $H$ consisting of some monochromatic $F$-copies, there are $O(n^{v(G)-1})$ non-rainbow copies of $G$.
If $H$ is rainbow $G$-free, then by statement ($\star$) and the hypergraph removal lemma, we may delete $o(n^{r})$ edges together with the corresponding monochromatic $F$-copies to obtain a $G$-free graph that still consists of at least $ex_{F}(n,G)-o(n^{r})$ edge-disjoint copies of $F$. This yields $ex_{F}(n,G)\leq\nu_{n}(F,G)+o(n^{r})$.
The statement ($\star$) clearly holds. Since every copy of $G$ contains two edges $e_{1},e_{2}$ of the same color (otherwise it would be rainbow) and $|e_{1}\cup e_{2}|\geq r+1$, there are at most
# $\mbox{ monochromatic }F\mbox{-copies}\cdot{e(F)\choose 2}=O(n^{r})$
choices of such monochromatic edges pairs $e_{1}$ and $e_{2}$, and hence there are $O(n^{r})\cdot n^{v(G)-(r+1)}=O(n^{v(G)-1})$ non-rainbow copies of $G$.
In what follows, we restrict our discussion to simple graphs. Imolay, Karl, Nagy and Váli [13] proved that $ex_{F}(n,G)=\Theta(n^{2})$ if and only if there is no homomorphism from $G$ to $F$, and for the nondegenerate case ($ex_{F}(n,G)=\Theta(n^{2})$),
$n^{2}/v(F)^{2}+o(n^{2})\leq ex_{F}(n,G)\leq ex(n,G)/e(F)+o(n^{2}).$ (1)
Li [15] extended the above results to uniform hypergraphs. We say $ex_{F}(n,G)$ attains its lower bound (resp. upper bound) in (1) if $ex_{F}(n,G)=n^{2}/v(F)^{2}+o(n^{2})$ (resp. $ex_{F}(n,G)=ex(n,G)/e(F)+o(n^{2})$). Note that $\chi(G)>\chi(F)$ implies that there is no homomorphism from $G$ to $F$. Imolay, Karl, Nagy and Váli [13] showed that $ex_{F}(n,G)$ attains its upper bound when $\chi(G)>\chi(F)$. In fact, using the stability of extremal graphs for $ex(n,G)$, one can show that this condition is also necessary: $ex_{F}(n,G)$ attains its upper bound if and only if $\chi(G)>\chi(F)$ [15].
Imolay, Karl, Nagy and Váli [13] posed the question of characterizing all graph pairs $(F,G)$ for which $ex_{F}(n,G)=n^{2}/v(F)^{2}+o(n^{2})$. The first such pair was found by Kovács and Nagy [14]: $F=C_{5}$ and $G=C_{3}$. Later, Balogh, Liebenau, Mattos and Morrison [1] refined this result by showing $(n^{2}-10n)/25\leq ex_{C_{5}}(n,C_{3})\leq(n^{2}+3n)/25+o(n)$. Characterizing when $ex_{F}(n,G)$ attains the lower bound is a challenging problem. Instead, we first study conditions that prevent it from attaining the lower bound.
Theorem 1.1.
For graphs $G$ and $F$ with $\chi(G)\leq\chi(F)$, if the following conditions hold, then $ex_{F}(n,G)$ does not attain its lower bound.
  1. (A).
    There are two nonadjacent vertices $u,v$ in $F$ such that $u,v$ have at most one common neighbor;
  2. (B).
    there is no homomorphism from $G$ to both $F$ and $F/\{u,v\}$.
Remark 1.1.
The condition (A) is necessary: for $t\geq 2$ and graph pair $F=C_{5}(t)$ and $G=C_{3}$, condition (B) holds for any nonadjacent $u,v$, yet $ex_{C_{5}(t)}(n,C_{3})$ attains the lower bound (see Theorem 1.4). Here $G(t)$ represents the $t$-blow-up of $G$.
The following result provides a method to generate more such pairs from one that attains the lower bound.
Theorem 1.2.
Let $G,F$ be two graphs with $ex_{F}(n,G)=\frac{n^{2}}{v(F)^{2}}+o(n^{2})$, and $s$ be a positive integer.
  1. 1.
    If $F(s)$ can be decomposed into $s^{2}$ edge-disjoint copies of $F$, then $ex_{F(s)}(n,G)=\frac{n^{2}}{v(F(s))^{2}}+o(n^{2})$.
  2. 2.
    Specifically, if $s$ is a prime with $v(F)\leq s$, then $ex_{F(s)}(n,G)=\frac{n^{2}}{v(F(s))^{2}}+o(n^{2})$.
For the degenerate case where $ex_{F}(n,G)=o(n^{2})$, we distinguish two subcases: the $(6,3)$-type bound $ex_{F}(n,G)=n^{2-o(1)}$ and other. Kovács and Nagy [14] generalized the $(6,3)$-theorem by showing that $ex_{F}(n,G)=n^{2-o(1)}$ if $G$ contains a triangle and there exists a homomorphism from $G$ to $F$. We explore graph pairs where $G$ is triangle-free and $ex_{F}(n,G)=n^{2-o(1)}$. The odd girth of $G$ is defined as the length of a shortest odd cycle in $G$. For each graph $G$, let
$\mathrm{Epi}(G)=\{G^{\prime}:\mbox{there is an epimorphism from }G\mbox{ to }G^{\prime}\}.$
Theorem 1.3.
  • (1).
    If $ex_{F}(n,G)=n^{2-o(1)}$ and no graph in $\mathrm{Epi}(G)$ with smaller order than $G$ is a subgraph of $F$, then for any positive integer $s$, we have $ex_{F(s)}(n,G)=n^{2-o(1)}$.
  • (2).
    If $g(F)=g(G)$, $\chi(F)=3$ and there exists a homomorphism from $G$ to $F$, then $ex_{C_{k}}(n,G)=n^{2-o(1)}$. Specifically, $ex_{C_{k}}(n,C_{k})=n^{2-o(1)}$.
Next, we study $ex_{F}(n,G)$ when $F$ and $G$ are cycles. Our investigation is motivated by the generalized Turán number $ex(n,F,G)$, which is the maximum number of copies of $F$ in an $n$-vertex $G$-free graph. The most famous problem in this area is a conjecture of Erdős [5] asserting that $ex(n,C_{5},C_{3})=\Theta(n^{5})$ (in fact, the extremal construction gives $(n/5)^{5}+o(n^{5})$). This conjecture was solved independently by Grzesik [11] and by Hatami, Hladký, Král, Norine, and Razborov [12] using flag algebras. Grzesik and Kielak [11] generalized the result by showing that $ex(n,C_{2r+1},C_{2r-1})=(n/(2r+1))^{2r+1}+o(n^{2r+1})$, and Beke and Janzer [2] characterized the extremal constructions. Bollobás and Győri [3] considered the opposite problem $ex(n,C_{3},C_{5})$, and Gishboliner and Shapira [9] studied further cases of “cycles versus cycles”. For a comprehensive survey on generalized Turán problems, we refer the reader to [8].
For cycles $F$ and $G$, we first extend the known results $F=C_{5},G=C_{3}$ to more cycle pairs that attain the lower bound.
Theorem 1.4.
If $k$ is odd and $\ell=k-2$, then for any positive integer $t$, $ex_{C_{k}(t)}(n,C_{\ell})=\frac{n^{2}}{(kt)^{2}}+o(n^{2})$.
If there exists a homomorphism from $C_{k}$ to $C_{\ell}$, then either $k$ is even, or both $k$ and $\ell$ are odd with $\ell\geq k$. Recall that $ex_{C_{3}}(n,C_{\ell})$ denotes the $3$-uniform linear Turán number of Berge-$C_{\ell}$. Ergemlidze, Győri and Methuku [6] proved that $ex_{C_{3}}(n,C_{4})=n^{3/2}/6+O(n)$ and $ex_{C_{3}}(n,C_{5})=n^{3/2}/(3\sqrt{3})+O(n)$; Füredi and Özkahya [7] established the general bound $ex_{C_{3}}(n,C_{2k+1})\leq 2kn^{1+1/k}+9kn$; Collier-Cartaino, Graber, and Jiang [4] extended the result to $ex_{C_{3}}(n,C_{k})\leq O(n^{1+1/\lfloor k/2\rfloor})$. We obtain the following general result.
Theorem 1.5.
Let $k,\ell\geq 1$ and $s\geq 3$ three two integers, and $F,G$ be two simple graphs.
  • (1).
    If $G$ is a bipartite graph, then $ex_{F}(n,G)=O(ex(n,G))$. Specifically, $ex_{C_{s}}(n,C_{2\ell})\leq O(n^{1+1/\ell})$.
  • (2).
    If $\ell>k$, then $ex_{C_{2k+1}}(n,C_{2\ell+1})\leq 8(k+1)^{2}\ell n^{1+1/(\ell-k+1)}$.
Remark 1.2.
Recall that for simple graphs $G$ and $F$, $ex_{F}(n,G)$ attains the upper bound if and only if $\chi(G)>\chi(F)$. Combining with Theorem 1.1, we have that $ex_{C_{k}}(n,C_{\ell})$ attains the upper bound in Ineq.(1) if and only if $k$ is even and $\ell$ is odd. Theorems 1.1 and 1.4 indicate that $ex_{C_{k}}(n,C_{\ell})$ attains the lower bound of Ineq.(1) if and only if $k,\ell$ are odd and $\ell=k-2$. Theorems 1.3 and 1.5 indicate that $ex_{C_{k}}(n,C_{\ell})=n^{2-o(1)}$ if and only if $k=\ell\geq 3$ are odd integers.
In the following result, we strengthen the first result in Theorem 1.5 when $\ell=4$.
Theorem 1.6.
  • (1).
    $\frac{\sqrt{2}n^{3/2}}{8}+o(n^{3/2})=ex(n/2,C_{4})\leq{ex}_{C_{4}}(n,C_{4})=\frac{\sqrt{2}n^{3/2}+n}{8}$.
  • (2).
    For $k\geq 5$, we have
[…]
Remark 1.3.
According to Theorem 1.2 in [16], the gap between the lower and upper bounds of $ex_{C_{4}}(n,C_{4})$ can be as large as $\Theta(n)$ for a positive density of integers $n$.
The paper is organized as follows. In section 2, we will prove some preliminary results; in Section 3, we prove Theorems 1.1 and 1.2; in Second 4, we prove Theorems 1.3, 1.4 and 1.5; In Second 5, we prove Theorem 1.6.
2 Preliminaries
For two graphs $G$ and $H$, let $G\choose H$} denote the set of all copies of $H$ in $G$. A map $\psi^{*}:{G\choose H}\rightarrow[0,1]$ such that for each $e\in E(G)$
$\sum_{e\in E(H^{\prime}),H^{\prime}\in{G\choose H}}\psi^{*}(H^{\prime})\leq 1$
is called a fractional $H$-packing of $G$. The maximum weight of $\sum_{H^{\prime}\in{G\choose H}}\psi^{*}(H^{\prime})$, denoted by $\nu^{*}_{G}(H)$, is called the fractional $H$-packing number of $G$. Recall that the $H$-packing number of $G$, denoted by $\nu_{G}(H)$, is the maximum number of edge-disjoint copies of $H$ in $G$. It is clear that $\nu^{*}_{G}(H)\geq\nu_{G}(H)$.
Theorem 2.1 (Haxell, Rödl).
For any fixed graph $H$ and an $n$-vertex graph $G$, $\nu^{*}_{G}(H)-\nu_{G}(H)=o(n^{2})$.
Lemma 2.1.
Suppose $H$ is an $n$-vertex graph consisting of monochromatic $F$-copies and $H$ is rainbow $G$-free, then we can remove at most $o(n^{2})$ monochromatic $F$-copies such that the resulting graph is $\mathrm{Epi}(G)$-free.
Proof.
By the statement ($\star$), there are $O(n^{v(G)-1})$ copies of $G$ in $H$. For each integer $1\leq\ell\leq v(G)-1$ and each $G^{\prime}\in\mathrm{Epi}(G)$ with $v(G^{\prime})=\ell$, there are $O(n^{\ell})$ copies of $G^{\prime}$. Consequently, by graph removal lemma, for each $G^{\prime\prime}\in\mathrm{Epi}(G)$, we can delete at most $o(n^{2})$ edges together with the corresponding monochromatic $F$-copies such that the resulting graph is $G^{\prime\prime}$-free. Since $|\mathrm{Epi}(G)|$ is finite, we can delete $o(n^{2})$ monochromatic $F$-copies such that the resulting graph is $\mathrm{Epi}(G)$-free. ∎
Lemma 2.2.
If $F$ is a graph of order $k$ and $G$ is an $n$-vertex graph consisting of $t$ monochromatic $F$-copies $F_{1},F_{2},\ldots F_{t}$, then there is a partition $V_{1},V_{2},\ldots,V_{k}$ of $V(G)$ and a bijection $f:[k]\rightarrow V(F)$, such that there are at least $t/k^{k}$ $F_{i}$s satisfying
  • (1).
    $|V(F_{i})\cap V_{j}|=1$ for each $j\in[k]$ (say $w_{j}=V(F_{i})\cap V_{j}$), and
  • (2).
    $w_{a}w_{b}\in E(F_{i})$ if and only if $f(a)f(b)\in E(F)$.
Proof.
Partition $V(G)$ into $V_{1},V_{2},\ldots,V_{k}$ randomly and uniformly (i.e. put each vertex of $H$ into each $V_{i}$ with the probability $1/k$). Let $\mathcal{S}=\{F_{j}:|V(F_{j})\cap V_{i}|=1\mbox{ for each }i\in[k]\}$. Since the expectation of $|\mathcal{S}|$ is $k!t/k^{k}$, there exists such a partition $V_{1},V_{2},\ldots,V_{k}$ of $V(G)$ such that $|\mathcal{S}|\geq k!t/k^{k}$. Since there are $k!$ bijection from $[k]$ to $V(F)$, there exists a bijection $f$ such that there are at least $t/k^{k}$ $F_{i}$s satisfying (2). ∎
The theta graph $\theta_{k}$ is a cycle $C_{k}$ with chord. Let $\theta_{\geq k}$ denote the set of all $\theta_{t}$ with $t\geq k$. In the following, we give a result by Füredi and Özkahya (see Lemma 7 in [7]).
Lemma 2.3 (Füredi and Özkahya).
Let $G$ be a $C_{2k+1}$-free graph and let $T$ be a subtree in $G$ with root $x$. Let $V_{i}$ denote the set of vertices of distance $i$ from $x$ in the tree $T$. Then for each $i\leq k$, $G[V_{i}]$ is $\theta_{\geq 2k}$-free.
3 Proofs of Theorems 1.1 and 1.2
In this section, we will prove Theorems 1.1 and 1.2.
Proof of Theorem 1.1: Since there is no homomorphism from $G$ to $F$, it follows that $\chi(G)\geq 3$. Therefore, $3\leq\chi(G)\leq\chi(F)$.
Let $v(F)=f$ and $V(F)=\{v_{1},v_{2},\ldots,v_{f-2},u,v\}$. Assume that $F^{\prime}=F/\{u,v\}$ is obtained by shrinking $u,v$ to the vertex $v_{0}$. Then $V(F^{\prime})=\{v_{1},v_{2},\ldots,v_{f-2},v_{0}\}$. If the common neighbor of $u,v$ exists, we can assume it as $v_{1}$.
Claim 1.
$e(F^{\prime}-\{v_{0},v_{1}\})\geq 1$.
Proof.
Suppose for the contrary that $e(F^{\prime}-\{v_{0},v_{1}\})=0$. Then $\{v_{2},v_{3},\ldots,v_{f-2}\},\{u,v\}$ and $\{v_{1}\}$ are three independent sets of $F$. Combining with $3\leq\chi(G)\leq\chi(F)$, we have that $\chi(F)=\chi(G)=3$. Since $v_{1}$ is the unique possible common vertex of $u$ and $v$, it follows that for each $2\leq i\leq f-2$, $v_{i}$ joins at most one vertex of $u,v$. Since $\chi(F)=3$, it follows that there exists a such integer $i$ such that either $v_{i}v_{1},v_{i}u\in E(F)$ or $v_{i}v_{1},v_{i}v\in E(F)$, implying $F$ contains a triangle. However, $\chi(G)=3$ indicates that there exists a homomorphism from $G$ to $F$, a contradiction. ∎
For sufficiently large $n$ and an $n$-vertex set $V$, let $\mathcal{P}=\{V_{0},V_{1},V_{2},\ldots,V_{f-2}\}$ be a partition of $V$ such that
$|V_{2}|=|V_{3}|=\ldots=|V_{f-2}|=\left\lfloor\frac{n}{2\sqrt{2}+f-3}\right\rfloor$
and $1\leq|V_{0}|-|V_{1}|\leq 1$. It is clear that
$|V_{0}|,|V_{1}|\geq\frac{\sqrt{2}n}{2\sqrt{2}+f-3}.$
Let $H$ be an $n$-vertex graph on $V$ with
$E(H)=\bigcup_{v_{i}v_{j}\in E(F^{\prime})}\{xy:x\in V_{i},y\in V_{j}\}.$
Since there is no homomorphism from $G$ to $F^{\prime}$, it follows that $H$ is $G$-free. By Claim 1, we can choose $V_{p},V_{q}\in\mathcal{P}$ such that $v_{p}v_{q}\in E(F^{\prime}-\{v_{0},v_{1}\})$. It is clear that
$e[V_{p},V_{q}]=|V_{p}||V_{q}|=\left\lfloor\frac{n}{2\sqrt{2}+f-3}\right\rfloor^{2}=\min\{e[V_{i},V_{j}]:v_{i}v_{j}\in E(F^{\prime})\}.$
Choose a positive real number
$\epsilon<\frac{3-2\sqrt{2}}{2\sqrt{2}-3+f}<\frac{1.8}{f-1.8}.$
Then $1/(2\sqrt{2}+f-3)>(1+\epsilon)/f$ and hence
$e[V_{p},V_{q}]>\frac{(1+\epsilon)^{2}n^{2}}{f^{2}}.$
Let $\mathcal{A}$ be the set of all copies $F_{0}$ of $F$ in $H$ such that $|V(F_{0})\cap V_{i}|=1$ for each $i\in[f-2]$ and $|V(F_{0})\cap V_{0}|=2$. For each edge $e\in E(H)$, let $\mathcal{A}_{e}$ denote the set of copies of $F$ in $\mathcal{A}$ containing $e$. By symmetry, for any two edges $e,e^{\prime}$ connecting the same parts in $\mathcal{P}$, $|\mathcal{A}_{e}|=|\mathcal{A}_{e^{\prime}}|$. Therefore, for $v_{i}v_{j}\in E(F^{\prime})$ and each $e\in E[V_{i},V_{j}]$, if $\{i,j\}\neq\{0,1\}$ or $\{i,j\}=\{0,1\}$ but $v_{1}$ is not the common neighbor of $u$ and $v$, then each $F_{0}\in\mathcal{A}$ contains exactly one edge in $E[V_{i},V_{j}]$, and hence $|\mathcal{A}_{e}|=|\mathcal{A}|/e[V_{i},V_{j}]$; if $\{i,j\}=\{0,1\}$ and $v_{1}$ is the common neighbor of $u$ and $v$, then each $F_{0}\in\mathcal{A}$ contains exactly two edges in $E[V_{i},V_{j}]$, and hence $|\mathcal{A}_{e}|=2|\mathcal{A}|/e[V_{i},V_{j}]$.
For each $F_{0}\notin\mathcal{A}$, define $\psi^{*}(F_{0})=0$; for each $F_{0}\in\mathcal{A}$, choose an edge $e^{*}\in E[V_{p},V_{q}]$ (note that $\{p,q\}\neq\{0,1\}$) and define $\psi^{*}(F_{0})=1/|\mathcal{A}_{e^{*}}|$, imlying $\psi^{*}(F_{0})=e[V_{p},V_{q}]/|\mathcal{A}|$. Note that for each edge $e\in E[V_{i},V_{j}]$, if $\{i,j\}\neq\{0,1\}$ or $\{i,j\}=\{0,1\}$ but $v_{1}$ is not the common neighbor of $u$ and $v$, then
$\sum_{F_{0}\in\mathcal{A}_{e}}\psi^{*}(F_{0})=|\mathcal{A}_{e}|\cdot e[V_{p},V_{q}]/|\mathcal{A}|=e[V_{p},V_{q}]/e[V_{i},V_{j}]\leq 1;$
if $\{i,j\}=\{0,1\}$ and $v_{1}$ is the common neighbor of $u$ and $v$, then
$\sum_{F_{0}\in\mathcal{A}_{e}}\psi^{*}(F_{0})=|\mathcal{A}_{e}|\cdot\frac{e[V_{p},V_{q}]}{|\mathcal{A}|}=\frac{2e[V_{p},V_{q}]}{e[V_{i},V_{j}]}\leq 2\left\lfloor\frac{n}{2\sqrt{2}+f-3}\right\rfloor^{2}\cdot\left(\frac{\sqrt{2}n}{2\sqrt{2}+f-3}\right)^{-2}\leq 1.$
Therefore, $\psi^{*}$ is a fractional $F$-packing of $H$, and
$\nu^{*}_{H}(F)\geq\sum_{F_{0}\in\mathcal{A}}\psi(F_{0})=\sum_{e\in E[V_{p},V_{q}]}\sum_{F_{0}\in\mathcal{A}_{e}}e[V_{p},V_{q}]/|\mathcal{A}|=e[V_{p},V_{q}]>\frac{(1+\epsilon)^{2}n^{2}}{f^{2}}.$
By Theorem 2.1, we have that
$\nu_{H}(F)\geq\nu^{*}_{H}(F)+o(n^{2})\geq\frac{(1+\epsilon)^{2}n^{2}}{f^{2}}+o(n^{2}).$
Consequently,
$ex_{F}(n,G)\geq\nu_{H}(F)\geq\frac{(1+\epsilon)^{2}n^{2}}{f^{2}}+o(n^{2}),$
as $H$ is $G$-free. $\square$
Proof of Theorem 1.2: We show the first statement. Suppose for the contrary that there exists $\epsilon>0$ such that $ex_{F(s)}(n,G)\geq\frac{n^{2}}{v(F(s))^{2}}+\epsilon n^{2}$ for infinitely many $n$. For each such $n$, let $H_{n}$ be a rainbow $G$-free graph consisting of $ex_{F(s)}(n,G)$ monochromatic $F(s)$-copies. By ($\star$) and graph removal lemma, we can delete at most $o(n^{2})$ monochromatic $F(s)$-copies from $H_{n}$ to obtain a $G$-free graph $H_{n}^{\prime}$. It is clear that $H^{\prime}_{n}$ consists of at least $\frac{n^{2}}{v(F(s))^{2}}+\epsilon n^{2}/2$ monochromatic $F(s)$-copies. Since each monochromatic $F(s)$-copy can be decomposed into $s^{2}$ edge-disjoint copies of $F$, $H_{n}^{\prime}$ can be decomposed into $s^{2}\cdot ex_{F(s)}(n,G)=\frac{n^{2}}{v(F)^{2}}+\epsilon s^{2}n^{2}/2$ edge-disjoint copies of $F$ (regard each of them as a monochromatic $F$-copies). Since $H^{\prime}_{n}$ is $G$-free, we have that $ex_{F}(n,G)\geq\frac{n^{2}}{v(F)^{2}}+\epsilon s^{2}n^{2}/2$ for infinitely many $n$, contradicting $ex_{F}(n,G)=\frac{n^{2}}{v(F)^{2}}+o(n^{2})$.
Now we show the second statement. For any prime $s$ with $v(F)\leq s$, it suffices to show that $F(s)$ can be decomposed into $s^{2}$ edge-disjoint copies of $F$. For convenience, let $v(F)=[t]$. Let $R$ be an $s$-blow up of $K_{t}$ such that the $t$ parts are $V_{1}=[s],V_{2}=[s],\ldots,V_{t}=[s]$. For any two vertices $x\in V_{1},y\in V_{2}$, define $K^{x,y}$ as a copy of $K_{t}$ induced by $x\in V_{1},y\in V_{2}$ and $x+iy\in V_{i+2}$ for each $i\in[t-2]$ (all values are taken modulo $s$).
Claim 2.
For two distinct pairs $(x,y)$ and $(x^{\prime},y^{\prime})$, where $x,x^{\prime}\in V_{1}$, $y,y^{\prime}\in V_{2}$, we have that $K^{x,y}$ and $K^{x^{\prime},y^{\prime}}$ are edge-disjoint.
Proof.
Suppose for the contrary that $e\in E(K^{x,y})\cap E(K^{x^{\prime},y^{\prime}})$. If $e$ is an edge between $V_{i+2},V_{j+2}$ for some $t-2\geq i>j\geq 1$, then $e=\{x+iy,x+jy\}=\{x^{\prime}+iy^{\prime},x^{\prime}+jy^{\prime}\}$. It is clear that $s|[(x+iy)-(x^{\prime}+iy^{\prime})]$ and $s|[(x+jy)-(x^{\prime}+jy^{\prime})]$. Hence, $s|(i-j)(y-y^{\prime})]$. Since $s$ is a prime, we have that either s|(ij)s|(i-j) or $s|(y-y^{\prime})$. Note that $0<i-j<t\leq s$ implies that $s$ does not divide iji-j. Consequently, $s|(y-y^{\prime})$. Since $y,y^{\prime}\in[s]$, it follows that $y=y^{\prime}$. Note that $s|[(x+iy)-(x^{\prime}+iy^{\prime})]$ implies that $s|(x-x^{\prime})$. Since $x,x^{\prime}\in[s]$, it follows that $x=x^{\prime}$, contradicting $(x,y)$ and $(x^{\prime},y^{\prime})$ are distinct pairs.
Now we assume that either $e=\{x,x+iy\}=\{x^{\prime},x^{\prime}+iy^{\prime}\}$ or $e=\{y,x+iy\}=\{y^{\prime},x^{\prime}+iy^{\prime}\}$ for some $i\in[t-2]$ (by symmetry, say $e=\{x,x+iy\}=\{x^{\prime},x^{\prime}+iy^{\prime}\}$). It is clear that $s|(x-x^{\prime})$ and $s|i(y-y^{\prime})$. Since $s$ is a prime, $0<i\leq s-2$ and $x,x^{\prime},y,y^{\prime}\in[s]$, it follows that $x=x^{\prime}$ and $y=y^{\prime}$, a contradiction. ∎
By Claim 2, $R$ can be decomposed into $s^{2}$ edge-disjoint copies of $K_{t}$. Note that $R-\bigcup_{ij\notin E(F)}E(V_{i},V_{j})$ is a $F(s)$ and $F^{x,y}=K^{x,y}-\bigcup_{ij\notin E(F)}E(V_{i},V_{j})$ is a copy of $F$ for each $x\in V_{1},y\in V_{2}$. Hence, $\{F^{x,y}:x\in V_{1},y\in V_{2}\}$ is an $F$-decomposition of $F(s)$. $\square$
4 Multicolor Turán number of cycles
Proof of Theorem 1.3: We show the first statement. Assume that $L=\lfloor n/s\rfloor$ and $H$ is an $L$-vertex rainbow $G$-free graph consisting of $ex_{F}(L,G)$ monochromatic $F$-copies. Then $ex_{F}(L,G)=L^{2-o(1)}$. For the $s$-blow-up $H(s)$ of $H$ (say each $v\in V(H)$ is replaced by $X_{v}$), color each edge $ab\in E(H(s))$ with $a\in X_{u}$ and $b\in X_{v}$ by the same color as the edge $uv$ in $H$. Then $H(s)$ is decomposed into $ex_{F}(L,G)=L^{2-o(1)}=n^{2-o(1)}$ edge-disjoint $F(s)$. It suffices to show $H(s)$ is $G$-free. Otherwise, let $G^{*}$ is a rainbow copy of $G$ in $H(s)$. For each $u\in V(G^{*})$, denote by $u^{\prime}$ be the unique vertex in $V(H)$ with $u\in X_{u^{\prime}}$. Let $S=\{u^{\prime}:u\in V(G^{*})\}$. If $|U|=v(G)$, then $H[U]$ contains a rainbow copy of $G$, a contradiction; if $|U|<v(G)$, then $H[U]$ contains a subgraph of order $|U|$ belonging to $\mathrm{Epi}(G)$, a contradiction. Therefore, $H(s)$ is $G$-free.
Now we show the second statement. Assume that $g(G)=g(F)=k$, where $k$ is odd. Since $\chi(F)=3$, there exists an integer $t$ such that $F$ are subgraphs of $C_{k}(t)$. Note that $F$ and $G$ contain $C_{k}$ as a subgraph. It is obvious that
$ex_{C_{k}(t)}(n,C_{k})\leq ex_{C_{k}(t)}(n,G)\leq ex_{F}(n,G).$
Since there exists a homomorphism from $G$ to $F$, $ex_{F}(n,G)=o(n^{2})$. It is sufficient to show that $ex_{C_{k}(t)}(n,C_{k})\geq n^{2-o(1)}$. By the first statement, we only need to show that $ex_{C_{k}}(n,C_{k})\geq n^{2-o(1)}$.
The proof depends on a known result by Ruzsa (see Theorem 2.3 in [17]), which states that there exists an $A\subseteq[n]$ of size $n^{1-O(\sqrt{\log n})}=n^{1-o(1)}$ such that $A$ does not contain a set of not all equal integers $x_{1},x_{2},\ldots,x_{k}$ satisfying $\sum_{i\in[k-1]}x_{i}=(k-1)x_{k}$. Let $n={k+1\choose 2}N$. We construct an $n$-vertex graph $H$ as follows: $V(H)$ can be partitioned into $k$ subsets $V_{1}=[N],V_{2}=[2N],\ldots,V_{k}=[kN]$. Edges in $H$ can appear only between parts $V_{i},V_{i+1}$ for each $i\in[k-1]$, and parts $V_{1},V_{k}$. For any vertices $x_{1}\in V_{1},x_{2}\in V_{2},\ldots,x_{k}\in V_{k}$, $x_{i}x_{i+1}\in E(H)$ if and only if $x_{i+1}-x_{i}\in A$ for each $i\in[k-1]$, and $x_{1}x_{k}\in E(H)$ if and only if $(x_{k}-x_{1})/(k-1)\in A$. It is clear that any $k$-cycle in $H$ intersects each $V_{i}$ in exactly one vertex, and each vertex of $V_{1}$ is contained in $|A|$ $k$-cycles. Hence, $H$ contains at least $|V_{1}||A|=n^{2-o(1)}$ $k$-cycles.
For a $k$-cycle $x_{1}x_{2}\ldots x_{k}x_{1}$ in $H$, it is clear that $c_{1}=x_{2}-x_{1},c_{2}=x_{3}-x_{2},\ldots,c_{k-1}=x_{k}-x_{k-1},c_{k}=(x_{k}-x_{1})/(k-1)$ are integers in $A$. However, $\sum_{i\in[k-1]}c_{i}=(k-1)c_{k}$, implying $c_{1}=c_{2}=\ldots=c_{k}$. Therefore, each edge between $V_{1}$ and $V_{2}$ of $H$ is contained in exactly one $k$-cycle. Hence, $H$ contains $n^{2-o(1)}$ edge-disjoint $k$-cycles (regard each such $k$-cycle as a monochromatic cycle of a unique color). Since each $k$-cycle in $H$ must contain an edge between $V_{1}$ and $V_{2}$, there is no rainbow $k$-cycle in $H$. Hence, $ex_{C_{k}}(n,C_{k})=n^{2-o(1)}$. $\square$
Proof of Theorem 1.4: We first show that $ex_{C_{k}}(n,C_{k-2})=n^{2}/k^{2}+o(n^{2})$. Assume that $G$ is a rainbow $C_{k-2}$-free graph consisting of $ex_{C_{k}}(n,C_{k-2})$ monochromatic $C_{k}$-copies. Since $\{C_{\ell}:3\leq\ell\leq k-2\mbox{ and }\ell\mbox{ is odd }\}\subseteq\mathrm{Epi}(C_{k})$. By Lemma 2.1, we can get an $n$-vertex graph $G^{\prime}$ consisting of $ex_{C_{k}}(n,C_{k-2})-o(n^{2})$ monochromatic $C_{k}$-copies such that $G^{\prime}$ is $C_{\ell}$-free for each odd $3\leq\ell\leq k-2$. Denote by $\mathcal{F}$ the set of monochromatic $C_{k}$-copies in $G^{\prime}$.
For each vertex $\in V(G), vv$ is contained in exactly $d_{G^{\prime}}(v)/2$ members in $\mathcal{F}$. Hence,
$\sum_{F\in\mathcal{F}}\sum_{v\in V(F)}d_{G^{\prime}}(v)=\sum_{v\in V(G)}\frac{d_{G^{\prime}}(v)^{2}}{2}\geq\frac{2e(G^{\prime})^{2}}{n}=\frac{2(k|\mathcal{F}|)^{2}}{n}.$ (2)
On the other hand, we have that for each $F\in\mathcal{F}$ and each vertex $v\in V(G)$, $|N_{G^{\prime}}(v)\cap V(F)|\leq 2$; otherwise, there is an odd cycle $C_{\ell}$ with $3\leq\ell\leq k-2$, a contradiction. Hence, for each $F\in\mathcal{F}$,
$\sum_{v\in V(F)}d_{G^{\prime}}(v)\leq 2v(G^{\prime})=2n.$
Thus, $\sum_{F\in\mathcal{F}}\sum_{v\in V(F)}d_{G^{\prime}}(v)\leq 2n|\mathcal{F}|$. Combining with Ineq. (2), we have that $|\mathcal{F}|\leq\frac{n^{2}}{k^{2}}$. Hence, $ex_{C_{k}}(n,C_{k-2})=\frac{n^{2}}{k^{2}}+o(n^{2})$.
Now we show that $ex_{C_{k}(t)}(n,C_{k-2})=n^{2}/k^{2}+o(n^{2})$. By the first statement of Theorem 1.2, we only need to show that $H=C_{k}(t)$ can be decomposed into $s^{2}$ edge-disjoint copies of $C_{k}$. Let $V_{1}=[t],V_{2}=[t],\ldots,V_{k}=[t]$ be the $k$ parts of $C_{k}(t)$. Edges of $H$ can only appear in $E[V_{1},V_{k}]$ or $E[V_{i},V_{i+1}]$ for $i\in[k-1]$. For each $x\in V_{1},y\in V_{2}$, let $C^{x,y}$ be the $k$-cycle with vertes set $\{x,y,x+y,x+2y,\ldots,x+(k-2)y,x\}$, where $x+iy\in V_{i+2}$. It suffices to show that for different edges $xy$ and $x^{\prime}y^{\prime}$, where $x,x^{\prime}\in V_{1}$ and $y,y^{\prime}\in V_{2}$, $C^{x,y}$ and $C^{x^{\prime},y^{\prime}}$ are edge-disjoint. Suppose for the contrary that $e$ is a common edge of $C^{x,y}$ and $C^{x^{\prime},y^{\prime}}$. If $e=\{x+iy,x+(i+1)y\}=\{x^{\prime}+iy^{\prime},x^{\prime}+(i+1)y^{\prime}\}$ for some $i\in[k-3]$ (i.e., $x+iy,x^{\prime}+iy^{\prime}\in V_{i+2}$), then $x=x^{\prime}$ and $y=y^{\prime}$, a contradiction. Similarly, the remaining two cases $e=\{x,x+(k-2)y\}=\{x^{\prime},x^{\prime}+(k-2)y^{\prime}\}$ and $e=\{y,x+y\}=\{y^{\prime},x^{\prime}+y^{\prime}\}$ both indicate $x=x^{\prime}$ and $y=y^{\prime}$, a contradiction. Hence, $C^{x,y}$ and $C^{x^{\prime},y^{\prime}}$ are edge-disjoint. $\square$
Proof of Theorem 1.5: (1). Let $v(F)=k$. It suffices to show that $ex_{F}(n,G)\leq k^{k}\cdot ex(n,G)$. Suppose, to the contrary, that $ex_{F}(n,G)>k^{k}\cdot ex(n,G)$. Let $H$ be an $n$-vertex rainbow $G$-free graph consisting of $ex_{F}(n,G)$ monochromatic $F$-copies. By Lemma 2.2, we can partition $V(H)$ into $V_{1},V_{2},\ldots,V_{k}$ so that there are at least $t=ex(n,G)+1$ monochromatic $F$-copies, say $F^{1},F^{2},\ldots,F^{t}$, satisfying two statements in Lemma 2.2. Let $i,j\in[k]$ be two indexs for which $e(V_{i},V_{j})\neq\emptyset$ and $|V_{i}\cup V_{j}|$ is maximized. Then $|V_{i}\cup V_{j}|\geq\frac{\delta(F)n}{e(F)}$. Note that $H^{\prime}=H[V_{i}\cup V_{j}]$ is a rainbow graph and $e(H^{\prime})\geq t>ex(n,G)\geq ex(|V_{i}\cup V_{j}|,G)$. Hence, $H^{\prime}$ contains a $G$, contradicting the fact that $H$ is rainbow $G$-free.
(2). Let $G$ be an $n$-vertex rainbow $C_{2\ell+1}$-free graph consisting of $ex_{C_{2k+1}}(n,C_{2\ell+1})$ monochromatic $C_{2k+1}$-copies. Let $\delta$ denote the average number of monochromatic $C_{2k+1}$-copies containing a vertex. Then
$e(G)=(2k+1)ex_{C_{2k+1}}(n,C_{2\ell+1})=(2k+1)\delta n.$ (3)
Delete all vertices $v$ with $d_{G}(v)<2\delta$, together with all monochromatic $C_{2k+1}$-copies incident to $v$. In the resulting graph, if there exist vertices $w$ of degree less than $2\delta$ (i.e., vertices incident to fewer than $\delta$ monochromatic $C_{2k+1}$-copies), delete these vertices as well as all monochromatic $C_{2k+1}$-copies containing any of them. Repeating this process, we eventually obtain a graph $G^{*}$. Since $e(G^{*})>e(G)-(2k+1)\delta n>0$, $G^{*}$ is a nonempty graph and has minimum degree at least $2\delta$.
For two integer sequences $\alpha=(a_{1},a_{2},\ldots,a_{k})$ and $\beta=(b_{1},b_{2},\ldots,b_{k})$, we say that $\alpha$ is lexicographically larger than $\beta$ if there exists $i\in[k]$ such that $a_{i}>b_{i}$ and $a_{j}=b_{j}$ for all $j<i$. Choose a vertex $x$ in $G^{*}$ and let $T_{x}$ be a rainbow tree rooted at $x$ of height at most $\ell$ (i.e., the maximum distance from $x$ to any vertex in $T_{x}$ is at most $\ell$). Then $V(T_{x})$ can be partitioned into $\ell+1$ sets $\{x\},V_{1},V_{2},\ldots,V_{\ell}$, where $V_{i}$ denotes the set of vertices at distance $i$ from $x$ in $T_{x}$. Note that if the height of $T_{x}$ is $i<\ell$, then $V_{j}=\emptyset$ for each $j>i$. Define $\gamma(T_{x})=(|V_{1}|,|V_{2}|,\ldots,|V_{\ell}|)$. Among all rainbow trees in $G^{*}$ rooted at $x$ of height at most $\ell$, select one, denoted $T$, that maximizes $\gamma(T)$ with respect to the lexicographic order.
Let $U_{i}$ denote the set of vertices at distance $i$ from $x$ in the rainbow tree $T$, and let $\mathcal{C}_{all}$ denote the set of monochromatic $C_{2\ell+1}$-copies that are incident to some vertex in $\bigcup_{0\leq i\leq\ell}U_{i}$. We partition $\mathcal{C}_{all}$ into two subsets
$\mathcal{T}=\{C\in\mathcal{C}_{all}:E(C)\cap E(T)\neq\emptyset\}\mbox{ and }\mathcal{C}=\{C\in\mathcal{C}_{all}:E(C)\cap E(T)=\emptyset\}.$
Define a subset $\mathcal{C}^{*}\subseteq\mathcal{C}$ by
$\mathcal{C}^{*}=\{C\in\mathcal{C}:V(C)\cap U_{i}\neq\emptyset\mbox{ for some }0\leq i\leq\ell-k\}.$
Let $G_{1}$ and $G_{2}$ be two subgraphs of $G^{*}$ induced by edge-sets $\bigcup_{C\in\mathcal{T}}E(C)$ and $\bigcup_{C\in\mathcal{C}^{*}}E(C)$, respectively.
Claim 3.
For each $C\in\mathcal{C}^{*}$,
  • (1).
    every edge in $E(C)$ lies either in some $U_{i}$ or between consecutive parts $U_{i}$ and $U_{i+1}$, and
  • (2).
    $V(C)\subseteq V(T)$.
Proof.
We show the first statement. Suppose, to the contrary, that there exists an edge $e=ab$ in $E(C)$ such that $a\in U_{i}$ and $b\in U_{j}$ for some $i,j\in[\ell]$ with $j-i\geq 2$. Then $T\cup e$ contains a unique cycle $Q$, and $Q$ contains an edge $f$ that lies between $U_{j}$ and $U_{j-1}$. Let $T^{\prime}=T+e-f$. Since the colors of $e$ and every edge in $T$ are distinct, $T^{\prime}$ is a rainbow tree rooted at $x$ of height at most $\ell$. However, since $j-i\geq 2$, we have $\gamma(T^{\prime})>\gamma(T)$, a contradiction.
Now we show the second statement. Suppose, to the contrary, that $V(C)-V(T)\neq\emptyset$. Let $i$ be the minimum integer such that $V(C)\cap U_{i}\neq\emptyset$. Since $C\in\mathcal{C}^{*}$, it follows that $i\in[\ell-k]$. Define
$Z_{p}=\{w:dist_{C}(w,V(C)\cap U_{i})=p\}.$
Since $V(C)-V(T)\neq\emptyset$, there exists a minimum integer $p\geq 1$ such that $Z_{p}-V(T)\neq\emptyset$ (say $w^{\prime}\in Z_{p}$). Then there exists an edge $e=w^{\prime}w^{\prime\prime}$ in $C$ with $w^{\prime\prime}\in Z_{p-1}$. By the first statement, we have $Z_{p-1}\subseteq U_{i+p-1}$. Since $i\in[\ell-k]$ and $p\leq k$, it follows that $i+p-1\leq\ell-1$. Hence, $T^{\prime}=T\cup e$ is a rainbow tree rooted at $x$ of height no more than $\ell$. Clearly, $\gamma(T^{\prime})>\gamma(T)$, again a contradiction. ∎
We further partition $\mathcal{T}$ into $\ell$ parts: for $i\in[\ell]$, let
$\mathcal{T}_{i}=\{C\in\mathcal{T}:E(C)\cap E(T)\mbox{ is an edge between }U_{i-1}\mbox{ and }U_{i}\}.$
Clearly, for each $i\in\ell$,
$|\mathcal{T}_{i}|=|U_{i}|.$ (4)
For each $C\in\mathcal{C}^{*}$, let
$top_{C}=\min\{i:U_{i}\cap V(C)\neq\emptyset\}.$
Observe that, by the maximality of $\gamma(T)$, there are no monochromatic $C_{2k+1}$-copies in $\mathcal{C}^{*}$ that intersect $x=U_{0}$. Hence, we can further partition $\mathcal{C}^{*}$ into $\ell-k$ subsets: for each $i\in[\ell-k]$, let $\mathcal{C}_{i}=\{C\in\mathcal{C}^{*}:top_{C}=i\}$.
Claim 4.
Assume that $C$ is a monochromatic $C_{2k+1}$-copy containing some vertex in $U_{i}$. If $C\in\mathcal{T}$, then $C\in\bigcup_{i\in[\max\{0,i-k\},i+k]}\mathcal{T}_{j}$; if $C\in\mathcal{C}^{*}$, then $C\in\bigcup_{i\in[\max\{0,i-k\},i]}\mathcal{C}_{j}$.
Proof.
If $C\in\mathcal{C}^{*}$, then the result follows immediately from statement (1) in Claim 3. Hence, assume that $C\in\mathcal{T}$. Without loss of generality, let $p$ be the integer such that $C\in\mathcal{T}_{p}$. By the definition of $\mathcal{T}_{p}$, there exists an edge of $C$, say $e=ab$, that also belongs to $T$, with $a\in U{p-1}$ and $b\in U_{p}$.
Suppose for the contrary that $p\notin[\max\{0,i-k\},i+k]$. Then |pi|>k|p-i|>k. Since $C$ contains a vertex in $U_{i}$ (say $z$) and $dist_{C}(z,b)\leq k$, there exists an integer $q\geq i$ and an integer $\alpha\geq 2$ such that $q+\alpha\leq p$ and $C$ contains an edge $f=a^{\prime}b^{\prime}$ with $a^{\prime}\in U_{q}$ and $b^{\prime}\in U_{q+\alpha}$. Clearly,
$T^{\prime}=T\left[\bigcup\nolimits_{0\leq j\leq q+1}U_{j}\right]\cup\{f\}$
is a new rainbow tree rooted at $x$ of height at most $\ell$, and $\gamma(T^{\prime})>\gamma(T)$, a contradiction. ∎
Since each $C\in\mathcal{C}^{*}$ is an odd cycle, by statement (1) in Claim 3, there exists an $i\in[\ell]$ and an edge of $C$, denoted by $e_{C}$, such that $e_{C}\in G^{*}[U_{i}]$. Let $R=\{e_{C}:C\in\mathcal{C}\}$ and $R_{i}=R\cap G^{*}[U_{i}]$. Clearly, each $E(T)\cup R_{i}$ is rainbow. Since $G^{*}$ is rainbow $C_{2\ell+1}$-free, by Lemma 2.3, $G^{*}[R_{i}]$ is $\theta_{\geq 2\ell}$-free. Since $ex(N,\theta_{\geq 2\ell})\leq 2(\ell-1)N$ (which is used in [7]), it follows that
$|R_{i}|=e(G^{*}[R_{i}])\leq 2(\ell-1)|U_{i}|.$
Let $I_{p}=[\max\{0,p-k\},p+k]$. It is obvious that for each $i\in[\ell-k]$, $I_{p}\subseteq[0,\ell]$.
Claim 5.
For each $i\leq\ell-k$, $\sum_{j\in[\max\{0,i-k\},i]}|\mathcal{C}_{j}|\leq 2(\ell-1)\sum_{j\in I_{i}}|U_{j}|$.
Proof.
Given an $i\in[\ell-k]$, statement (1) in Claim 3 implies that for each $C\in\mathcal{C}_{i}$, $V(C)\subseteq\bigcup_{i\leq j\leq i+k}U_{j}$. Hence, $e_{C}$ lies in some $U_{j}$, $i\leq j\leq i+k$. Consequently,
$\sum_{j\in[\max\{0,i-k\},i]}|\mathcal{C}_{j}|\leq\sum_{j\in I_{i}}|R_{j}|\leq 2(\ell-1)\sum_{j\in I_{i}}|U_{j}|.$
Note that for each monochromatic $C_{2k+1}$-copy $C$, if $V(C)\cap U_{i}\neq\emptyset$, then $C$ contributes at most
$\omega_{i}(C)\leq 2e(C)=4k+2$
in $\sum_{v\in U_{i}}d_{G^{*}}(v)$. By Claim 4, $\mathcal{T}$ contributes
[…]
to $\sum_{v\in U_{i}}d_{G^{*}}(v)$, and $\mathcal{C}^{*}$ contribute
$\sum_{j\in[\max\{0,i-k\},i]}\sum_{C\in\mathcal{C}_{j}}\omega_{i}(C)\leq(4k+2)\sum_{j\in[\max\{0,i-k\},i]}|\mathcal{C}_{j}|\leq 2(\ell-1)(4k+2)\sum_{j\in I_{i}}|U_{j}|$
to $\sum_{v\in U_{i}}d_{G^{*}}(v)$ (the last inequality holds due to Claim 5). Since each vertex in $U_{i}$, $0\leq i\leq\ell-k$, is incident only to monochromatic $C_{2k+1}$-copies belonging to $\mathcal{T}$ or $\mathcal{C}^{*}$, it follows that
$\sum_{v\in U_{i}}d_{G^{*}}(v)$ $\leq(4k+2)(2\ell-1)\sum_{j\in I_{i}}|U_{j}|.$
Since each $i$ is contained at most $2k+1$ intervals $I_{j}$, it follows that for each $t\in[\ell-k]$,
$\sum_{0\leq i\leq t}\sum_{j\in I_{i}}|U_{j}|\leq(2k+1)\sum_{0\leq i\leq t+k}|U_{i}|.$
Consequently,
$\sum_{0\leq i\leq t}\sum_{v\in U_{i}}d_{G^{*}}(v)$ $\leq(4k+2)(2\ell-1)\sum_{0\leq i\leq t}\sum_{j\in I_{i}}|U_{j}|$
$\leq 2(2k+1)^{2}(2\ell-1)\sum_{0\leq i\leq t+k}|U_{i}|.$
Recall that the minimum degree of $G^{*}$ is at least $2\delta$. Hence,
$2\delta\sum_{0\leq i\leq t}|U_{i}|\leq\sum_{0\leq i\leq t}\sum_{v\in U_{i}}d_{G^{*}}(v)\leq 2(2k+1)^{2}(2\ell-1)\sum_{0\leq i\leq i+k}|U_{i}|.$
Let $W_{i}=\bigcup_{0\leq j\leq i}U_{j}$. Then
$\delta|W_{t}|\leq 2(2k+1)^{2}(2\ell-1)\sum_{0\leq i\leq i+k}|U_{i}|\leq 8(k+1)^{2}\ell|W_{t+k}|.$
By the maximality of $\gamma(T)$, $|W_{1}|=|U_{1}|+1>\delta$. Hence,
$n\geq|W_{\ell}|\geq|W_{1}|\left(\frac{\delta}{8(k+1)^{2}\ell}\right)^{\ell-k}\geq\frac{\delta^{\ell-k+1}}{[8(k+1)^{2}\ell]^{\ell-k}},$
implying $\delta\leq 8(k+1)^{2}\ell n^{1/(\ell-k+1)}$. Since $ex_{C_{2k+1}}(n,C_{2\ell+1})=\delta n$ (see Eq. (3)), it follows that
$ex_{C_{2k+1}}(n,C_{2\ell+1})\leq 8(k+1)^{2}\ell n^{1+1/(\ell-k+1)},$
the proof thus completes. $\square$
5 Proof of Theorem 1.6
To establish the foundation for the subsequent proof, we first define a rainbow common neighbor and introduce a classification of vertex pairs based on this notion. A vertex $w$ is called a rainbow common neighbor of two distinct vertices $x$ and $y$ if $w$ is adjacent to both $x$ and $y$, and the colors on the edges $wx$ and $wy$ are distinct. Based on the number of such neighbors, we classify an unordered pair of distinct vertices $\{u,v\}$ as follows: If $\{u,v\}$ has three rainbow common neighbors, it is called a large vertex pair; if $\{u,v\}$ has two rainbow common neighbors, it is called a medium vertex pair; if $\{u,v\}$ has at most rainbow common neighbor, it is called a small vertex pair.
For a edge-colored graph $G$ and an edge $e\in E(G)$, we use $c(e)$ to denote the color of edge $e$, and use $C(G)$ to denote the set of colors appear in $G$. For two edge-colored graphs $G_{1}$ and $G_{2}$, we say $G_{1}$ and $G_{2}$ have the same coloring pattern if
  • $G_{1}$ is isomorphic to $G_{2}$, and
  • there exists an isomorphism $\phi:G_{1}\rightarrow G_{2}$ and a bijection $f:C(G_{1})\rightarrow C(G_{2})$ such that for any edge $uv$ of $G_{1}$, $f(c(uv))=c(\phi(u)\phi(v))$.
Figure 1: Two types of coloring patterns of $B^{u,v}$.
Lemma 5.1.
Let $G$ be a rainbow $C_{4}$-free graph consisting of monochromatic $C_{k}$-copies. Then for any two vertices $u,v$ of $G$,
  • if $k\geq 6$, then $u,v$ have at most six rainbow common neighbors; if $k=4,5$, then $u,v$ have at most three rainbow common neighbors.
  • if $k=5$, then the coloring pattern of $B^{u,v}$ is Type 1 or 2 (see Figure 1).
  • if $k=4$, then the coloring pattern of $B^{u,v}$ is Type 1 (see Figure 1).
Proof.
Suppose $u$ and $v$ share $d$ rainbow common neighbors, denoted $x_{1},x_{2},\ldots,x_{d}$. Without loss of generality, assume that $c(ux_{1})$ is black and $c(vx_{1})$ is red.
Since $G$ is rainbow $C_{4}$-free, for each $i\in\{2,3,\ldots,d\}$, we have $\{c(ux_{i}),c(vx_{i})\}\cap\{\text{red,black}\}\neq\emptyset$. Given that each of $u$ and $v$ is incident with exactly four edges colored either black or red, and $ux_{1}$, $vx_{1}$ account for two of these, it follows that $d\leq 7$.
Suppose, for contradiction, that $d=7$. Then all rainbow common neighbors appear as depicted in Figure 2(1). In this case, $c(ux_{4})=c(ux_{5})$ must be a new color, say blue. By the same reasoning, $c(ux_{3})$ is also blue. Consequently, $u$ is incident with three blue edges, a contradiction. Hence, $d\leq 6$.
Now assume $k=4,5$ and $d\geq 3$. Define two stars $S_{u}=G[ux_{1},\ldots,ux_{d}]$ and $S_{v}=G[vx_{1},\ldots,vx_{d}]$.
Case 1. $S_{u}$ or $S_{v}$ is not rainbow.
By symmetry, we may assume $S_{u}$ is not rainbow and that $c(ux_{2})=$ black. We claim $c(vx_{p})=\text{red}$ for each $3\leq p\leq d$. Indeed, if $c(vx_{p})=$ black, then $k=5$ and the black $C_{5}$ is $x_{p}x_{1}ux_{2}vx_{p}$, contradicting that $x_{2}$ is a rainbow common neighbor of $u$ and $v$; if $c(vx_{p})$ is neither black nor red (say blue), then since $ux_{1}vx_{p}u$ is not a rainbow $C_{4}$, we must have $c(ux_{p})=$red. Since $uv_{2}vx_{p}u$ is not a rainbow $C_{4}$, it follows that $c(vx_{2})$ is blue. Consequently, for each $3\leq p\leq d$, $c(vx_{2})=c(x_{p})=\text{blue}$, implying $d=3$. Moreover, the coloring pattern of $B^{u,v}$ is Type 1.
For the edges $ux_{3}$ and $vx_{2}$, it is clear that $\{c(ux_{3}),c(vx_{2})\}\cap\{\text{red},\text{black}\}=\emptyset$ (for otherwise, whenever $c(vx_{2})=\text{red}$ or $c(vx_{3})=\text{black}$, the black $C_{5}$ and the red $C_{5}$ have a common edge, a contradiction). Moreover, we have $c(vx_{2})=c(ux_{3})$, which must be a new color (say green); otherwise $ux_{3}vx_{2}u$ would form a rainbow $C_{4}$, a contradiction. Consequently, the coloring pattern of $B^{u,v}$ is Type 1.
Case 2. Both $S_{u}$ and $S_{v}$ are rainbow.
Without loss of generality, let $c(ux_{2})=$ red and $c(vx_{3})=$ black. Since both $S_{u}$ and $S_{v}$ are rainbow, $\{c(ux_{3}),c(vx_{2})\}\cap\{\text{red},\text{black}\}=\emptyset$ and $d=3$.
Since $ux_{3}vx_{2}u$ is not a rainbow $C_{4}$ and $\{c(ux_{3}),c(vx_{2})\}\cap\{\text{red},\text{black}\}=\emptyset$, we must have $c(ux_{3})=c(vx_{2})$, which is a new color (say green). Thus, for $k=4,5$, the coloring pattern of $B^{u,v}$ is Type 2. However, when $k=4$, the coloring pattern Type 2 would imply that $uv$ belongs to both a green and a red $C_{5}$, a contradiction. ∎
Figure 2: Rainbow common neighbors of $u,v$ (left side). The right side figure demonstrates the sharpness for the case $k=6$.
Lemma 5.2.
If $G$ is a rainbow $C_{4}$-free graph consisting of monochromatic $C_{4}$-copies, then
  • for any large vertex pair $\{u,v\}$, two of the the three pairs in their rainbow common neighbors are small vertex pairs, and
  • for any small vertex pair $\{x,y\}$, there are at most one large vertex pair whose rainbow common neighborhoods contain $\{x,y\}$.
Proof.
Let $\{u,v\}$ be a large vertex pair and $x_{1},x_{2},x_{3}$ be their rainbow common neighbors. Then the coloring pattern is of Type 1. Furthermore, there exists additional two vertices $y_{1},y_{2}$ such that the union of black, green and red $C_{5}$s is depicted in Figure 3.
Figure 3: The union of three monochromatic $C_{4}$s.
We complete the first statement by showing $\{x_{1},x_{3}\}$ and $\{x_{1},x_{2}\}$ are small vertex pairs. By symmetry, we only need to show that $\{x_{1},x_{3}\}$ is small vertex pair. Note that $u$ is a rainbow common neighbor of $\{x_{1},x_{3}\}$. Assume for the contrary that $z\neq u$ is also a rainbow common neighbor of $\{x_{1},x_{3}\}$. Since $zx_{1}ux_{3}z$ is not a rainbow $C_{4}$, it follows that $\{c(zx_{1}),c(zx_{3})\}\cap\{\text{black,green}\}\neq\emptyset$. Hence, $z\in\{x_{2},y_{1}\}$. If $z=x_{2}$, then $x_{2}x_{1}\in E(G)$ and $c(x_{2}x_{1})\notin\{\text{red,blue,green}\}$. However, $x_{1}x_{2}uvx_{1}$ is a rainbow $C_{4}$, a contradiction. Similarly, If $z=y_{1}$, then $y_{1}x_{2}vx_{3}y_{1}$ is a rainbow $C_{4}$, a contradiction.
Next, we show the second statement. Suppose to the contrary that $\{x,y\}$ is a small vertex pair that is contained in the rainbow common neighborhoods of two different large vertex pairs $\{u,v\}$ and $\{u^{\prime},v^{\prime}\}$. Assume that the coloring pattern of $B^{u,v}$ is shown in Figure 1(1). Then either $\{x,y\}=\{x_{1},x_{2}\}$ or $\{x,y\}=\{x_{1},x_{3}\}$ (by symmetry, say $x=x_{1}$ and $y=x_{2}$). If $\{u,v\}\cap\{u^{\prime},v^{\prime}\}=\emptyset$, then one of $\{u,v\}$ and one of $\{u^{\prime},v^{\prime}\}$ are rainbow common neighbors of $x_{1},x_{2}$, a contradiction. Hence, $\{u,v\}\cap\{u^{\prime},v^{\prime}\}\neq\emptyset$ (without loss of generality, let $v=v^{\prime}$).
Since $x_{1}$ and $x_{2}$ are rainbow common neighbors of $u^{\prime}$ and $v$, we have ${\text{green},\text{red}}\subseteq C(B^{u^{\prime},v})$. By the characterization of coloring pattern Type 1, $u^{\prime}$ lies in either the red or the green $C_{5}$. Hence $u^{\prime}\in\{x_{3},y_{2}\}$. In either case, one of $x_{1}$ or $x_{2}$ is not a rainbow common neighbor of $u^{\prime}$ and $v$, a contradiction. ∎
Proof of Theorem 1.6: For $k\geq 5$, let $G$ be an $n$-vertex rainbow $C_{4}$-free graph consisting of $ex_{C_{k}}(n,C_{4})$ monochromatic $C_{k}$-copies. Let
[…]
Then for any two vertices $u,v$ of $G$, there are at most $d(k)$ rainbow common neighbors in $G$. Hence,
# $\mbox{ rainbow }K_{1,2}\leq d(k){n\choose 2}$.
On the other hand, since each $v\in V(G)$ belongs to exactly $d_{G}(v)/2$ monochromatic $C_{k}$s,
# $\mbox{ Rainbow }K_{1,2}=4\sum_{v\in G}{d(v)/2\choose 2}=\sum_{v\in V(G)}\left(\frac{d(v)^{2}}{2}-d(v)\right)\geq\frac{1}{2n}\left(\sum_{v\in G}d(v)\right)^{2}-2e=\frac{2e^{2}}{n}-2e.$
Hence, $\frac{2e^{2}}{n}-2e\leq d(k){n\choose 2}$. If $k\geq 6$, then $e(G)\leq(\sqrt{6}n^{3/2}+n)/2$, and hence
$ex_{C_{k}}(n,C_{4})=e(G)/k\leq\frac{\sqrt{6}n^{3/2}+n}{2k}.$
If $k=5$, then $e(G)\leq(\sqrt{3}n^{3/2}+n)/2$, and hence
$ex_{C_{5}}(n,C_{4})=e(G)/5\leq\frac{\sqrt{3}n^{3/2}+n}{10}.$
For $k=4$, we will give a more subtle discussion. Let $\mathcal{P}$ be the set of $(P_{s},P_{\ell})$, where $P_{s}$ is a small vertex pair and $P_{\ell}$ is a large vertex pair whose rainbow common neighborhood contains both vertices in $P_{s}$. Assume that there are $s$ small vertex pairs, $m$ middle vertex pairs and $\ell$ large vertex pairs in $G$. Then $s+m+\ell={n\choose 2}$. The first statement of Lemma 5.2 implies $|\mathcal{P}|=2\ell$, and the second statement of Lemma 5.2 implies $|\mathcal{P}|\leq s$. Hence, $\ell\leq s/2$. Since each large vertex pair has three rainbow common neighbors, each middle vertex pair has two rainbow common neighbors and each small vertex pair has at most one rainbow common neighbors, it follows that
# $\mbox{ rainbow }K_{1,2}\leq 3\ell+2m+s=2{n\choose 2}+\ell-s\leq 2{n\choose 2}$.
Consequently, $2{n\choose 2}\geq\frac{2e^{2}}{n}-2e$, implying $e(G)\leq(n+\sqrt{2}n^{3/2})/2$ and
$ex_{C_{4}}(n,C_{4})=e(G)/4\leq\frac{n+\sqrt{2}n^{3/2}}{8}.$
We now construct a lower bound. Let $G$ be an $n$-vertex $C_{4}$-free graph with
$e(G)={ex}(n,C_{4})=\frac{1}{2}n^{\frac{3}{2}}+o(n^{\frac{3}{2}}).$
Define $G^{\prime}$ as the $2$-blow-up $G(2)$ of $G$. Then $|G^{\prime}|=2n$. In this construction, each edge of $G$ is replaced by a $C_{4}$ in $G^{\prime}$, and each vertex $v\in V(G)$ is replaced by a $2$-set, denoted by $X_{v}$. Clearly, $G^{\prime}$ consists of
$ex(n,C_{4})=\frac{1}{2}n^{\frac{3}{2}}+o(n^{\frac{3}{2}})=\frac{\sqrt{2}v(G^{\prime})^{3/2}}{8}+o(v(G^{\prime})^{3/2})$
monochromatic $C_{4}$-copies. Assign a distinct color to each of these $C_{4}$’s. Next, we show that $G^{\prime}$ contains no rainbow $C_{4}$. Suppose, to the contrary, that $G^{\prime}$ contains a rainbow $C_{4}$, and denote it by $a_{1}a_{2}a_{3}a_{4}a_{1}$. Then there exist vertices $v_{1},v_{2},v_{3},v_{4}$ in $G$ such that $v_{i}\in X_{a_{i}}$ for each $i\in[4]$. It is easy to verify that $v_{1},v_{2},v_{3},v_{4}$ are distinct, and hence $v_{1}v_{2}v_{3}v_{4}v_{1}$ is a rainbow $C_{4}$ in $G$, a contradiction. Therefore,
$ex_{C_{4}}(n,C_{4})\geq ex(n/2,C_{4})\geq\frac{\sqrt{2}n^{3/2}}{8}+o(n^{3/2}).$
$\square$
6 Acknowledgments
Ping Li is Supported by the National Natural Science Foundation of China No. 12201375. We are grateful to Suyun Jiang for recommending reference [2], which is helpful for the proof of Theorem 1.4.